OpenStudy (anonymous):

Solve the proportion: 8/x=3/9

OpenStudy (unklerhaukus):

first step is to simplify the fraction on the right hand side, can you cancel a common factor?

OpenStudy (anonymous):

whats the answer

OpenStudy (texaschic101):

your not even gonna try to work it out ?

OpenStudy (anonymous):

i got 24

OpenStudy (texaschic101):

thats correct

OpenStudy (anonymous):

What is the solution of the equation 3x − 7 = 4x?

OpenStudy (texaschic101):

first subtract 3x from both sides

OpenStudy (anonymous):

7

OpenStudy (texaschic101):

not 7.....-7 3x - 7 = 4x -3x -3x ------------ -7 = x

OpenStudy (anonymous):

What is the solution of the equation x − 8 = 17?

OpenStudy (texaschic101):

add 8 to both sides

OpenStudy (anonymous):

9

OpenStudy (texaschic101):

x - 8 = 17 -- add 8 to both sides +8 +8 ------------ x = 25